# MiLiC-Eval: Benchmarking Multilingual LLMs for China’s Minority Languages

### v1

@@ 4.2 Faithful Evaluation of LRL Abilities
Prior studies on LRLs often rely on a single task format for evaluation, such as simple NLU tasks Yong et al. (2023); Luukkonen et al. (2023); Lin et al. (2024) or translation NLLB Team et al. (2024). However, such an over-reliance on a single task type may lead to biased assessments of model capabilities. Additionally, many studies use machine-translated data for evaluation Hu et al. (2020); Chen et al. (2024), but the inherent noise in such data can obscure the true performance of models. In contrast, MiLiC-Eval uses diverse task formats derived from the same set of documents and recruits native speakers to translate the data. We discuss how these designs can provide a more robust and faithful assessment of model performance in LRLs.
@@
We argue that reliance on a single evaluation task can result in a skewed assessment of a model’s true capabilities in LRLs. To demonstrate this, we select two task groups from MiLiC-Eval, where the tasks within each group are derived from the same set of texts. We evaluate these tasks across different sizes of Qwen-2.5. The results are presented in Figure 4.
As the overall capabilities of models increase with the scaling of sizes, the speed of improvement varies significantly across different task formats. NLU tasks such as topic classification improves rapidly while translation tasks show a much slower rate of improvement. This indicates that relying on simple NLU tasks might lead to over-optimistic conclusions while solely using translation tasks for evaluation might underestimate models’ abilities. In contrast, by diversifying task formats through task parallelism, MiLiC-Eval can provide a more faithful and comprehensive assessment of LLMs’ understanding of LRLs.
@@
To evaluate the effects of using MT systems for data construction, we use NLLB-200-3.3B NLLB Team et al. (2024), a widely-used MT model, to translate several tasks in MiLiC-Eval from Chinese into minority languages [^4] Note that despite its wide language coverage, NLLB-200 does not support the traditional Mongolian script.. Table 3 shows model performance on the machine-translated task data.
We observe a consistent performance drop with machine-translated data compared to human-translated data, particularly on the math reasoning task, which demands more accurate translation. The drop is more pronounced in languages with poor MT support, such as Tibetan. According to the report of NLLB-200, Tibetan receives the weakest support among the three languages, which might explain the largest performance decline.
We further check the translation results of NLLB0-200 and find that repetition errors are frequent, where the model keeps outputting the same words Wang et al. (2024b). Additionally, the model often omits or modifies numerical values when translating math problems. These findings underscore the importance of native speakers in collecting evaluation data for LRLs.
@@ 4.3 Skill-wise Tracking of Language Adaptation
Figure 5 presents the continual pretraining course of the two languages. For Uyghur, to which Qwen-2.5 has had some exposure during pretraining, we observe a consistent improvement across all evaluated abilities. In particular, lexicology and topic modeling show the most significant gains, whereas abilities such as syntax, generation, translation, and symbolic reasoning exhibit modest or little improvements.
In contrast, Qwen-2.5 has limited exposure to Mongolian, and its tokenizer exhibits poor support, tokenizing most Mongolian characters as bytes. Consequently, continual pretraining yields minimal improvements on Mongolian. However, after expanding the tokenizer by 3K tokens obtained from the pretraining corpus Hewitt (2021), we observe noticeable improvements in several abilities for Mongolian, especially in lexicology and topic modeling, similar to the patterns observed on Uyghur. This finding challenges previous studies claiming that vocabulary expansion has little effect on downstream performance Csaki et al. (2024). We also note an initial decline in performance for certain Mongolian abilities at the start of training after vocabulary expansion. This drop may be due to the model’s adaptation to the newly added vocabulary, after which performance improves as the model adjusts to the expanded lexicon.
@@ 6 Conclusion
We present MiLiC-Eval, a multilingual benchmark comprising 24K instances across 9 tasks and 4 minority languages in China. MiLiC-Eval is distinguished by three key features: (1) a focus on underrepresented writing systems, (2) fine-grained skill-wise evaluation, and (3) cross-language and cross-task parallelism. We hope that MiLiC-Eval will not only advance LLM support for LRLs in China but also inspire research on underrepresented languages in other regions, such as Africa, India, and Southeast Asia.
@@ Appendix B Details of Data Collection
Table 4: ISO codes, languages families, and writing systems of the languages in MiLiC-Eval.
@@ B.1 Collection of Topic Classification (Passage)
We then select four labels for each language, which are easy to distinguish. We show the selected labels of each language in Table 5. Because there are not enough qualified labels for Uyghur, we additionally collect articles from a Uyghur news website nur.cn. We obtain 150 instances from the Heath and Sport columns respectively.
@@ B.2 Human Annotation
Table 5: The labels and sources of the instances in topic classification (Passage)
@@ Appendix C Implementation Details
Model Variants Used in Evaluation
Several LLMs have multiple variants available. Here we report the variants used in our evaluation in Section 3. We use Llama-3.1-8B-Instruct for LLama-3.1, aya-expanse-8b for Aya-Expanse, gemma-2-9b-it for Gemma-2, Qwen2.5-7B-Instruct for Qwen-2.5, Ministral-8B-Instruct-2410 for Ministral, gpt-4o-mini-2024-07-18 for GPT-4o-mini, emma-500-llama2-7b for EMMA-500, bayling-2-llama-3-8b for BayLing-2, LLaMAX3-8B-Alpaca for LLaMAX-3, and TowerInstruct-Mistral-7B-v0.2 for TowerInstruct. Our use of these models is consistent with their intended use.
@@
For the tasks in MiLiC-Eval, we use 5-shot exemplars, except for Title Generation, for which we use 3-shot exemplars due to the long lengths of news articles.
@@ D.1 Full Evaluation Results on MiLiC-Eval
In Table 6 we report the scores of each model on each task in MiLiC-Eval. Additionally, we report the performance of Qwen-2.5-7B, Llama-3.1-8B, and Gemma-2-9B on the Chinese and English versions of the tasks.
@@ D.2 Effect of Model Sizes
Uyghur
@@
Kazakh
@@
Mongolian
@@
High-Resource Languages
Qwen-2.5-7B (zh) - 87.1 - 92.5 90.6 - 54.0 25.3 64.2 23.4 76.3 -
Llama-3.1-8B (zh) - 86.3 - 90.2 83.8 - 49.5 25.0 62.6 24.0 69.5 -
Gemma-2-9B (zh) - 88.5 - 94.2 92.2 - 56.8 31.2 69.4 25.9 64.3 -
Qwen-2.5-7B (en) - 86.6 - 90.5 86.3 - - - - - 71.2 -
Llama-3.1-8B (en) - 86.5 - 87.8 85.1 - - - - - 75.0 -
Gemma-2-9B (en) - 88.0 - 93.7 92.4 - - - - - 65.7 -
Table 6: Scores (%) of different LLMs on MiLiC-Eval. Vocab. Test refers to Vocabulary Understanding. TC Sent. refers to topic classification (Sentence). TC Pass. refers to topic classification (Passage). Read. Comp. refers to Reading Comprehension. Resp. Sel. refers to Response Selection. Title Gen. refers to Title Generation. MT-A refers to Machine Translation (Article). xx2en denotes translation from the minority languages to English. en2xx denotes translation from English to the minority languages. MT-D refers to Machine Translation (Dialogue). Math refers to Math Reasoning. When calculating the average, the score for the two translation tasks is the mean of xx2en and en2xx. The highest scores are indicated in bold, while the second-highest scores are marked with underline.

### v2

@@ 4.2 Faithful Evaluation of LRL Abilities
Prior studies on LRLs often rely on a single task format for evaluation, such as simple NLU tasks Yong et al. (2023); Luukkonen et al. (2023); Lin et al. (2024) or translation NLLB Team et al. (2024). However, such an over-reliance on a single task type may lead to biased assessments of model capabilities. Additionally, many studies use machine-translated data for evaluation Hu et al. (2020); Chen et al. (2024); Huang et al. (2025), but the inherent noise in such data can obscure the true performance of models. In contrast, MiLiC-Eval uses diverse task formats derived from the same set of documents and recruits native speakers to translate the data. We discuss how these designs can provide a more robust and faithful assessment of model performance in LRLs.
@@
We argue that reliance on a single evaluation task can result in a skewed assessment of a model’s true capabilities in LRLs. To demonstrate, we select two task groups from MiLiC-Eval, where the tasks within each group are derived from the same set of texts. We evaluate these tasks across different sizes of Qwen-2.5. The results are presented in Figure 5.
As the models’ overall capabilities increase with the scaling of sizes, the speed of improvement varies significantly across different task formats. NLU tasks, such as topic classification, improve rapidly, while translation tasks show a much slower rate of improvement. This indicates that evaluation on simple NLU tasks only might lead to over-optimistic conclusions, while solely using translation tasks for evaluation might underestimate models’ abilities. In contrast, by diversifying task formats through task parallelism, MiLiC-Eval can provide a more faithful and comprehensive assessment of LLMs’ understanding of LRLs.
@@
To evaluate the effect of using MT systems for data construction, we use NLLB-200-3.3B NLLB Team et al. (2024), a widely-used MT model, to translate several tasks in MiLiC-Eval from Chinese into minority languages [^4] Note that despite its wide language coverage, NLLB-200 does not support the traditional Mongolian script.. Table 3 shows the model performance on machine-translated task data.
We observe a consistent performance drop with machine-translated data compared to human-translated data, particularly on the math reasoning task, which demands a more accurate translation. The drop is more pronounced in languages with poor MT support, such as Tibetan. The chrF++ scores of translation results are 23.0 for Tibetan, 30.4 for Uyghur, and 26.3 for Kazakh. Tibetan has the lowest chrF++ score among the three languages, which also exhibits the most significant performance decline resulting from translated evaluation data.
We further check the translation results of NLLB0-200 and find that the most commonly observed error is repetition, where the model keeps outputting the same words Wang et al. (2024b). Additionally, the model often omits or modifies numerical values when translating math problems. These findings underscore the importance of native speakers in collecting evaluation data for LRLs.
@@ 4.3 Skill-wise Tracking of Language Adaptation
Figure 6 presents the continual pretraining course of the two languages. For Uyghur, to which Qwen-2.5 has had some exposure during pretraining, we observe a consistent improvement across all evaluated abilities. In particular, lexicography and topic modeling show the most significant gains, whereas abilities such as syntax, generation, translation, and symbolic reasoning exhibit modest or little improvement.
In contrast, Qwen-2.5 has limited exposure to Mongolian, and its tokenizer exhibits poor support, tokenizing most Mongolian characters as bytes. Consequently, continual pretraining yields minimal improvements on Mongolian. However, after expanding the tokenizer by 3K tokens obtained from the pretraining corpus Hewitt (2021), we observe noticeable improvement in several abilities for Mongolian, especially in lexicography and topic modeling, similar to the patterns observed on Uyghur. This finding challenges previous studies claiming that vocabulary expansion has little effect on downstream performance Csaki et al. (2024). We also note an initial decline in performance for certain Mongolian abilities at the beginning of training after vocabulary expansion. This drop may be due to the model’s adaptation to the newly added vocabulary, after which performance improves as the model adjusts to the expanded lexicon.
@@ 6 Conclusion
We present MiLiC-Eval, a multilingual benchmark comprising 24K instances across 9 tasks and 4 minority languages in China. MiLiC-Eval is distinguished by three key features: (1) a focus on underrepresented writing systems, (2) cross-language and cross-task parallelism, and (3) fine-grained skill-wise evaluation. We hope that MiLiC-Eval will not only advance LLM support for LRLs in China but also inspire research on underrepresented languages in other regions, such as Africa, India, and Southeast Asia.
@@
Acknowledgements
This work is supported in part by NSFC (62161160339) and Beijing Science and Technology Program (Z231100007423011). We thank the anonymous reviewers for their valuable suggestions. We also thank all the annotators who contributed to the data collection. For any correspondence, please contact Yansong Feng.
@@ Appendix B Details of Data Collection
Table 4: ISO codes, language families, and writing systems of the languages in MiLiC-Eval.
@@ B.1 Collection of Topic Classification (Passage)
We then select four labels for each language, which are easy to distinguish. We show the selected labels of each language in Table 5. Because there are not enough qualified labels for Uyghur, we additionally collect articles from a Uyghur news website nur.cn. We obtain 150 instances from the Health and Sport columns, respectively.
@@ B.2 Human Annotation
Table 5: The labels and sources of the instances in topic classification (Passage).
Series Used Checkpoints
Ministral Ministral-8B-Instruct-2410
Aya-Expanse aya-expanse-8b
Qwen-2.5 Qwen2.5-7B-Instruct
Qwen-3 Qwen3-8B, Qwen3-14B
Llama-3.1 Llama-3.1-8B-Instruct
Gemma-2 gemma-2-9b-it
Gemma-3 gemma-3-12b-it
TowerInstruct TowerInstruct-Mistral-7B-v0.2
EMMA-500 emma-500-llama2-7b
BayLing-2 bayling-2-llama-3-8b
LLaMAX-3 LLaMAX3-8B-Alpaca
GPT-4o gpt-4o-mini-2024-07-18
GPT-4.1 gpt-4.1-2025-04-14
Gemini-2.0 gemini-2.0-flash-001
Table 6: The model variants used in evaluation.
@@ Appendix C Implementation Details
Models Used in Evaluation
Several LLMs have multiple variants available. Here we report the variants used in our evaluation in Table 6.
@@
For the tasks in MiLiC-Eval, we use 5-shot examples, except for Title Generation, for which we use 3-shot examples due to the long lengths of news articles.
@@ D.1 Full Evaluation Results on MiLiC-Eval
In Table 7, Table 8, Table 9, and Table 10, we report the scores of each model on each task in MiLiC-Eval.
Additionally, we report the performance of Qwen-2.5-7B, Llama-3.1-8B, and Gemma-2-9B on the Chinese and English versions of the tasks in Table 2.
@@ D.2 Effect of Model Sizes
GPT-4o-mini 66.0 40.7 60.5 40.0 38.8 24.3 23.2 16.3 21.5 18.0 19.5 36.6
GPT-4.1 85.3 66.7 84.4 64.0 65.8 32.1 35.4 24.6 38.3 25.1 53.0 57.0
Gemini-2.0-Flash 91.7 84.5 92.3 87.0 89.4 42.1 48.8 33.4 53.8 34.2 84.0 72.9
Table 7: Scores (%) of different LLMs on the Tibetan tasks of MiLiC-Eval. Vocab. Test refers to Vocabulary Understanding. TC Sent. refers to topic classification (Sentence). TC Pass. refers to topic classification (Passage). Read. Comp. refers to Reading Comprehension. Resp. Sel. refers to Response Selection. Title Gen. refers to Title Generation. MT-A refers to Machine Translation (Article). xx2en denotes translation from the minority languages to English. en2xx denotes translation from English to the minority languages. MT-D refers to Machine Translation (Dialogue). Math refers to Math Reasoning. When calculating the average, the score for the two translation tasks is the mean of xx2en and en2xx. The highest scores are indicated in bold, while the second-highest scores are marked with underline.
Model Vocab. TC TC Read. Resp. Title MT-A MT-A MT-D MT-D Math Avg.
Test Sent. Pass. Comp. Sel. Gen. xx2en en2xx xx2en en2xx
Random 25.0 16.2 25.0 27.0 25.0 -   -   -   -   -   - -
@@
GPT-4o-mini 92.4 67.6 92.5 75.5 80.8 18.8 48.6 28.3 44.5 23.8 72.0 63.6
GPT-4.1 95.2 87.3 96.4 83.5 89.4 26.6 54.1 37.5 51.5 31.3 82.5 72.0
Gemini-2.0-Flash 97.2 85.0 94.6 88.5 92.1 32.0 56.2 48.7 55.2 39.2 86.0 75.0
Table 8: Scores (%) of different LLMs on the Uyghur tasks of MiLiC-Eval. Vocab. Test refers to Vocabulary Understanding. TC Sent. refers to topic classification (Sentence). TC Pass. refers to topic classification (Passage). Read. Comp. refers to Reading Comprehension. Resp. Sel. refers to Response Selection. Title Gen. refers to Title Generation. MT-A refers to Machine Translation (Article). xx2en denotes translation from the minority languages to English. en2xx denotes translation from English to the minority languages. MT-D refers to Machine Translation (Dialogue). Math refers to Math Reasoning. When calculating the average, the score for the two translation tasks is the mean of xx2en and en2xx. The highest scores are indicated in bold, while the second-highest scores are marked with underline.
Model Vocab. TC TC Read. Resp. Title MT-A MT-A MT-D MT-D Math Avg.
Test Sent. Pass. Comp. Sel. Gen. xx2en en2xx xx2en en2xx
Random 25.0 16.2 25.0 27.0 25.0 -   -   -   -   -   - -
@@
GPT-4o-mini 79.8 63.4 73.4 63.5 61.4   6.0 40.0   1.3 34.7   1.2 54.0 48.9
GPT-4.1 89.7 85.0 79.8 82.0 89.2 31.1 57.9 15.4 46.0 13.1 70.0 65.9
Gemini-2.0-Flash 95.9 84.5 77.2 90.0 91.1 33.9 63.0 36.3 49.9 24.3 78.5 70.9
Table 9: Scores (%) of different LLMs on the Kazakh tasks of MiLiC-Eval. Vocab. Test refers to Vocabulary Understanding. TC Sent. refers to topic classification (Sentence). TC Pass. refers to topic classification (Passage). Read. Comp. refers to Reading Comprehension. Resp. Sel. refers to Response Selection. Title Gen. refers to Title Generation. MT-A refers to Machine Translation (Article). xx2en denotes translation from the minority languages to English. en2xx denotes translation from English to the minority languages. MT-D refers to Machine Translation (Dialogue). Math refers to Math Reasoning. When calculating the average, the score for the two translation tasks is the mean of xx2en and en2xx. The highest scores are indicated in bold, while the second-highest scores are marked with underline.
Model Vocab. TC TC Read. Resp. Title MT-A MT-A MT-D MT-D Math Avg.
Test Sent. Pass. Comp. Sel. Gen. xx2en en2xx xx2en en2xx
Random 25.0 16.2 25.0 27.0 25.0 -   -   -   -   -   - -
@@
GPT-4o-mini 30.2 18.5 35.5 36.5 32.9   6.4 19.6   0.8 16.6   0.8   7.0 20.7
GPT-4.1 34.1 21.1 46.2 45.0 43.5   9.2 20.1   0.6 16.6   7.2 23.5 27.2
Gemini-2.0-Flash 81.2 82.4 93.1 84.5 86.5 21.8 52.7 12.1 48.1 20.1 85.5 66.8
Table 10: Scores (%) of different LLMs on the Mongolian tasks of MiLiC-Eval. Vocab. Test refers to Vocabulary Understanding. TC Sent. refers to topic classification (Sentence). TC Pass. refers to topic classification (Passage). Read. Comp. refers to Reading Comprehension. Resp. Sel. refers to Response Selection. Title Gen. refers to Title Generation. MT-A refers to Machine Translation (Article). xx2en denotes translation from the minority languages to English. en2xx denotes translation from English to the minority languages. MT-D refers to Machine Translation (Dialogue). Math refers to Math Reasoning. When calculating the average, the score for the two translation tasks is the mean of xx2en and en2xx. The highest scores are indicated in bold, while the second-highest scores are marked with underline.
Model TC Read. Resp. MT-A MT-A MT-D MT-D Math
Sent. Comp. Sel. xx2en en2xx xx2en en2xx
Tibetan
Qwen-2.5-7B 30.6 48.3 39.6 18.8 10.8 14.2 13.2 13.3
Llama-3.1-8B 63.5 44.8 37.4 27.5   9.3 28.3 12.8 29.7
Gemma-2-9B 61.8 58.2 53.0 25.2 15.6 27.8 19.6 32.7
Uyghur
Qwen-2.5-7B 70.6 57.0 47.8 30.8 12.2 26.4 11.7 38.0
Llama-3.1-8B 76.9 59.2 46.0 41.3 17.6 36.2 15.9 48.0
Gemma-2-9B 75.9 58.5 59.5 37.0 12.1 31.7 11.1 46.8
Kazakh
Qwen-2.5-7B 56.5 51.7 40.1 25.6   1.0 21.0   1.1 21.3
Llama-3.1-8B 55.5 45.7 32.9 24.3 11.0 18.9 11.0 24.0
Gemma-2-9B 57.9 51.7 43.0 23.4   7.8 19.9   8.7 29.3
Mongolian
Qwen-2.5-7B 30.1 44.0 40.4 18.1   4.8 14.1   4.0 11.2
Llama-3.1-8B 23.1 31.5 25.6 18.4   5.5 14.9   6.7   8.7
Gemma-2-9B 19.9 42.8 37.0 17.4   5.1 15.5   6.7 14.7
Chinese
Qwen-2.5-7B 87.1 92.5 90.6 54.0 25.3 64.2 23.4 76.3
Llama-3.1-8B 86.3 90.2 83.8 49.5 25.0 62.6 24.0 69.5
Gemma-2-9B 88.5 94.2 92.2 56.8 31.2 69.4 25.9 64.3
English
Qwen-2.5-7B 86.6 90.5 86.3 - - - - 71.2
Llama-3.1-8B 86.5 87.8 85.1 - - - - 75.0
Gemma-2-9B 88.0 93.7 92.4 - - - - 65.7
Table 11: Scores (%) of three LLMs on the tasks with cross-lingual parallism in MiLiC-Eval. TC Sent. refers to topic classification (Sentence). Read. Comp. refers to Reading Comprehension. Resp. Sel. refers to Response Selection. MT-A refers to Machine Translation (Article). xx2en denotes translation from the minority languages to English. en2xx denotes translation from English to the minority languages. MT-D refers to Machine Translation (Dialogue). Math refers to Math Reasoning. .